# hyperrefSuppressing link with empty target

### v1

@@ 4.2 Meridional Circulation
(a) 0.016 bar
(b) 0.2 bar
Figure 15: Temporally averaged zonal wind (arrows) and temperature profile (map) at two different pressures (0.0026 bar - left, and 0.2 bar - right) for a non-synchronous HD209458b-like model in which the difference between the rotational and orbital periods is large ($\Omega_{\mathrm{rot}}=4\Omega_{\mathrm{orb}}=4\Omega_{0}$). Here we have forgone showing the weakly non-synchronous regime as we have already established its similarity to its synchronous equivalent. Note that the x-axis of these plots is relative-longitude, indicating that we have taken the temporal mean of the wind/temperature profile in the frame of reference of the irradiation.
4.3 Horizontal-Wind and Temperature Profiles
@@
Briefly, when the difference between the orbital and rotational periods is small, we find a wind/temperature structure that is almost identical to that found in the classical hot Jupiter regime. This includes a strong equatorial jet which is driven by off-equator $m=1$ standing wave circulations, as well as the classical thermal butterfly. Note that we find slightly more asymmetry in the flows here than in the synchronous case, asymmetry which we attribute to the time-dependent radiative forcing profile at least slightly disrupting the $m=1$ standing wave structure. Other evidence for the time-dependent radiative forcing can be found in the temperature profile of the non-synchronous mean night-side, which a hint hotter here than in an equivalent synchronous model, likely as a result of remnant thermal energy (that is still dissipating) from when that physical region of the atmosphere was irradiated.
On the other hand, when the difference between the rotational and orbital periods is large, other than in the fully radiatively controlled regions of the outer atmosphere, significant changes to the wind and temperature structure become evident, changes which we can be seen in Figure 15. For example, at $0.016$\mathrm{bar}$$, we find strong thermal ‘wings’ that extend from mid-latitudes on the ‘mean’ (where the mean is taken in the frame of reference of the irradiation) day-side to near the equator on the ‘mean’ night-side, ‘wings’ which trace the relatively weak off-equator standing wave circulation the develops here. These ‘wings’ only increase in strength as we move deeper into the atmosphere ($0.2$\mathrm{bar}$$ 15(b)), where we find that the non-synchronous mean day-side temperature profile has becoming increasingly smeared eastwards of the (moving) sub-stellar point. Whilst this description make the temperature profile sound a little like the classical thermal butterfly, differences in advective transport mean that this is not the case.
@@
Note that the dynamics are slightly different when the rotational period is less than the orbital period. This occurs for two reasons, the first is that the direction of the hot-spot shift is reversed westwards rather than eastwards) and the second in the decreasing influence of rotation at low $\Omega_{\mathrm{rot}}$ resulting in divergent rather than rotational dynamics. Note that the aforementioned smearing of the temperature profile by the shifting sub-stellar point, as well as the above factors, means that we also no longer find any evidence for the formation of the night-side hot-spot that was a defining feature of our slow-rotation-regime models (subsection 3.3).
@@
Figure 16: Global mean vertical enthalpy flux for 13 non-synchronous HD209458b-like models spanning the rotation range $\Omega_{\mathrm{rot}}=0.5\Omega_{Orbit}\rightarrow 4\Omega_{Orbit}$, where $\Omega_{Orbit}=\Omega_{0}$ is held constant. Note that negative fluxes indicate net downward enthalpy transport.
@@ 5 Discussion and Conclusion
This analysis revealed that, for the rotation rate ranges considered, our models fall into one of a number of distinct regimes.
For synchronous rotation we find three distinct dynamical regimes: slow rotation when $\Omega_{\mathrm{rot}}<0.5\Omega_{0}$, the classical hot Jupiter regime when $\sim 0.5\Omega_{0}<\Omega_{\mathrm{rot}}<\sim 5\Omega_{0}$ (an analysis of which can be found in the main text), and the rapidly rotating regime when $\Omega_{\mathrm{rot}}\gg 5\Omega_{0}$.
@@
We find that the slowly rotating regime is characterised by a slower zonal-mean zonal-jet (1(a)) which does not extend deep into the atmosphere ($P>0.5$\mathrm{bar}$$ is quiescent). The underlying cause of this slow wind can be understood via the helmholtz wind decomposition, which reveals that the strongest component of the horizontal wind, in the outer atmosphere, is the divergent component associated with a global overturning circulation as heat rises on the day-side, diverges from the sub-stellar point, before converging on the night-side, just to the west of the sub-stellar point. Note that the exact location at which the winds converge is dependent upon the rotation rate, with the convergence point shifting westwards towards the eastern terminator as the influence of rotation grows. This divergent wind acts as the main driver of horizontal enthalpy transport, leading to the formation of a rather unexpected feature: a night-side hot-spot (5(g) and 10(a)). This night-side hot-spot appears to be a robust feature of the slowly rotating regime and plays a role in shaping the mid-atmosphere dynamics. For example, around $0.2$\mathrm{bar}$$, we find that the night-side hot spot is hotter than any point on the day-side, an effect which reshapes the winds, driving a strong night-day flow.
@@
As a result we find that the global overturning circulation, associated with the divergent day-side wind and a broad sub-stellar point upflow, is closed in the mid-atmosphere, helping to explain the quiescent deep atmosphere. This effect can be seen in the meridional circulation profile, where we find an equatorial downflow that is confined to low pressures, and horizontal mixing at the bottom of the jet, where the night-side hot-spot forms. Finally this lack of a strong equatorial downflow that connects the outer atmosphere with the deep, $P>10\,$\mathrm{bar}$$, atmosphere means that the net deep heating is also weak (11(a) and 11(d)), leading to a deep atmosphere that is not significantly heated (10(a)), thus suggesting that slowly rotating warm/hot Jupiters should exhibit little to no radius inflation.
@@
On the other hand, we find that the rapidly rotating regime is characterised by a narrow, but deep, equatorial jet braced by easterly counterflows (1(c)). This is a result of the increasing influence that the Coriolis effect plays on the atmospheric dynamics, suppressing high-latitude off-equator flows. Specifically we find that, as the rotation rate rises, the standing $m=1$ Rossby and Kelvin waves which drive the archetypal super-rotating jet become increasingly tilted and confined to low latitudes, leading not only to the narrower jet seen here as the circulation cells become latitudinally compressed, but also a weakening of the overall equatorial winds as the angular momentum pumping (from mid-latitudes to the equator) mechanism becomes less efficient. Note that this compression only grows as we increase the rotation rate, leading to noticeably slow jets at very high $\Omega_{\mathrm{rot}}$ (7(a)). This latitudinal compression of the zonal-winds and off-equator flows also impacts the advection of enthalpy away from the sub-stellar point. For example, we find that the thermal butterfly (5(c)/h/l) has become somewhat confined to equatorial regions, with eastwards advection at the equator, westwards/polewards advection at low latitudes and almost no advection at higher latitudes where the Coriolis forces are strongest (and winds are weakest).
@@
Vertically, the latitudinal compression and resulting banding of the zonal-mean zonal-wind leads to a meridional circulation profile (4(c)) which contains multiple circulation cells in each hemisphere, although these circulations cells do still combine at the equator to drive a weakened downflow at all pressures, connecting the outer and deep atmospheres. Note that, much like the zonal wind, the meridional circulation is significantly weakened at high latitudes thanks to the Coriolis effect, hence most vertical mixing is also confined to low-latitudes (i.e. the equatorial region). This has important implications for the vertical enthalpy transport and hence heating of the deep atmosphere, both of which we find to be globally weak (9(l), 11(f), and 10(c)). Note however that this is just the global value. If we instead consider the equatorial mean (where vertical mixing is strongest), we find a net downward enthalpy flux comparable to that found in the classical hot Jupiter regime, and which leads to slight warming of the equatorial regions deep T-P profile (and only the equatorial regions due to the quiescence of the off-equator deep atmospheric winds). As such, we might expect such objects to be slightly puffed up near the equator, although such an effect is likely to remain minor and may be entangled with other effects that rotation and irradiation has on the atmospheres shape.
@@
This may help to explain the radius inflation, or lack-thereof, observed for ultra-hot Jupiters: planets with an orbital period of less than around 20 hours tend to exhibit little to no radius inflation e.g. Wasp-18b - Arcangeli et al. 2019 - or Wasp-43b - Hellier et al. 2011), whereas planets with orbital periods of $>22$ hours start to show signs of inflation (e.g. Wasp-103b - Gillon et al. 2014 - or KELT-16b - Oberst et al. 2017), similar to the transition seen with rotation rate here, although so far our models have not identified such a tight transition. This is also in agreement with the results of Sainsbury-Martinez et al. (2021), who found similar dynamics and a similar lack of radius inflation for SDSS1411b.
@@
Thanks to JWST, as well as other next-generation ground and space based telescopes, there is likely to be in influx of long orbital period hot/warm Jupiters in the near future. Not only would these objects likely fall into the slowly rotating regime, it is also increasingly likely that at least some will not be perfectly synchronously rotating. This can occur because the synchronisation time-scale increases with orbital radii, approaching the system age for a hot/warm Jupiter orbiting a sun-like star at $0.2$\mathrm{ua}$$. There are also a number of scenarios which might lead to a non-synchronous but highly irradiated hot Jupiter being observed. For example, we know that hot Jupiters cannot have formed so close to their host star and hence must have undergone planetary migration (Lin et al. 1996; Heller 2019). If we where to observe a hot Jupiter that had recently, or is still undergoing, migration, then it might not yet have had time to synchronise.
@@
As a result, and since many of the dynamics observed in hot Jupiter atmospheres are directly or indirectly linked to the strong day/night thermal forcing (from the day/night divergent flows in the slowly rotating regime to the thermally driven, $m=1$ Kelvin and Rossby standing waves which are responsible for the equatorial jet when rotational effects dominate), we next investigated how varying levels of non-synchronicity, as characterised by a difference between the varying planetary rotation rate and a fixed orbital period affects atmospheric dynamics. Here we found two additional regimes, depending upon the difference between the planetary rotational and orbital periods: when the difference is small, on the order of a few percent ($<15\%$) of the smaller period, we find dynamics that are qualitatively similar to the synchronous regime, just a little weaker. On the other hand when the difference is large, we find significant differences in the dynamics, especially in the driving of the super-rotating jet.
@@
Briefly, when the difference between the rotational and orbital periods is small, we find dynamics, including vertical enthalpy transport, that are structurally similar to those found in the synchronous regime, but just a little bit weaker. The biggest difference comes about when we look at the horizontal temperature profile in the outer atmosphere, where we find that the ‘sweeping’ of the stellar irradiation essentially acts as a form of day-night transport, with residual heat leading to a warmer night-side and hence a reduced day-night temperature difference.
@@
The above holds true when we also vary $\Omega_{\mathrm{orb}}$. As such, we expect small difference between the rotational and orbital periods to have little effect on observable features, such as the phase curve (hot-spot location) or radius. Furthermore, since the synchronisation timescale rapidly decreases with decreasing orbital radii it is highly likely that the majority of warm/hot Jupiters detected in the near future will be either synchronously rotating or only weakly non-synchronous. Whilst this is good in one sense, i.e. we can understand the dynamics of such objects quite easily, it also means that, with current tools, there is a distinct possibility that some weakly non-synchronous objects (which are potentially young, depending upon the local synchronisation time-scale), might be erroneously labelled as tidally-locked.
@@
But what about a scenario in which a planet has recently migrated onto a shorter orbit and not yet synchronised its rotation with its host star, leading to a large difference between the rotational and orbital angular rotation rates? Here our models suggest that the dynamics should be distinct from those found in the synchronous regime. For example, we find a zonal-mean zonal-jet (12(b)) which is both slower and shallower than its synchronous counterpart. This has a knock-on effect on the vertical dynamics, weakening both the meridional circulation and vertical enthalpy flux, the latter of which rapidly decreases as the difference between the rotational and orbital periods grows (Figure 16). The underlying cause of these changes can be best understood by exploring the horizontal wind/temperature profiles in the outer atmosphere (Figure 15). One immediate observation is that the thermal butterfly that typically forms in the mid-atmosphere of non-slowly-rotating hot Jupiters is not present. Instead, we find that the temperature profile has become longitudinally smeared, with the strength of the smearing depending upon $\Delta\Omega$: i.e. as the difference between the rotational and orbital periods increases, the time since the night-side was last irradiated gets smaller, leading to increased residual heating and a hot-spot that appears to trail the irradiation. Not only does this have the effect of reducing the strength of the day/night forcing, it also prevents the formation of the standing-wave structures which drive much of the dynamics of hot Jupiters, including the super-rotating jet and hence meridional circulation and vertical enthalpy flux. This can also shift the transition between the rotational-dominated and divergence-dominated dynamical regimes to faster planetary rotation rates, which will lead to additional changes in the strength of the vertical advection, since divergence-dominated winds are less efficient at vertical transport.
@@
Overall, this suggests that hot/warm Jupiters with large differences between their rotational and orbital periods should be observationally distinct from a synchronously rotating planet at the same orbital radius. This includes both a reduced phase curve amplitude as well as a reduced hot-spot offset, the latter of which could be either to the east or the west of the sub-stellar point depending upon the sign of the difference between the rotational and orbital periods. We might also expect to see a reduced level of radius inflation for such objects thanks to the reduced vertical heat flux, although this conclusion is more tenuous since newly formed Jupiter-like objects which have recently undergone planetary migration are likely still dissipating excess energy, associated with for example gravitational collapse, from the formation process, meaning that their envelopes are likely puffed up (Szulágyi et al. 2016). Such a scenario could be confirmed by comparing observed radii with planetary formation models, such as those of Fortney et al. (2007).
@@
In this work we have discussed how rotation alone can significantly impact and alter the dynamics of Hot Jupiters, ranging from the outer atmosphere wind and temperature structure, to the vertical transport of enthalpy. The latter effect may be of particular importance when trying to explain variations in hot Jupiter radii found at fixed surface irradiation, in particular the lack of inflation observed for some very-short-orbit hot Jupiters and hot Brown Dwarfs.
@@
One interesting discovery which might have significant implications for atmospheric dynamics and chemistry is the formation of a night-side hot-spot (via divergent day-night winds) in the mid-atmosphere of our slowly and synchronously rotating regime models. If such a night-side hot-spot is a robust feature of more slowly rotating hot/warm Jupiters, something that we wish to investigate using a next-generation GCM (such as expeRT/MITgcm) that robustly models radiative heating and cooling at all pressure levels, the associated temperature inversion may significantly alter atmospheric chemistry by acting, for example, as a cold trap and hence reshaping the vertical distribution of key observable models. Understanding such an effect would likely require modelling the atmospheres of slowly rotating hot/warm Jupiters with a GCM that includes both robust radiative transport and dis-equilibrium chemistry, something that is currently far too computationally expensive to robustly investigate.

### v2

@@ 4.2 Meridional Circulation
(a) 0.0026 bar
(b) 0.016 bar
(c) 0.2 bar
Figure 16: Temporally averaged, over five Earth-years of simulation time, zonal wind (arrows) and temperature maps at three different pressures ( 0.0026 bar - left, 0.016 bar - centre, and 0.2 bar - right) for a non-synchronous HD209458b-like model in which the difference between the rotational and orbital periods is large ($\Omega_{\mathrm{rot}}=4\Omega_{\mathrm{orb}}=4\Omega_{0}$). Here we have forgone showing the weakly non-synchronous regime as we have already established its similarity to its synchronous equivalent. Note that the x-axis of these plots is relative-longitude, indicating that we have taken the temporal mean of the wind/temperature in the frame of reference of the irradiation.
4.3 Horizontal-Wind and Temperature Maps
@@
Briefly, when the difference between the orbital and rotational periods is small, we find a wind/temperature structure that is almost identical to that found in the classical hot Jupiter regime. This includes a strong equatorial jet which is driven by off-equator $m=1$ standing wave circulations, as well as the classical thermal butterfly. Note that we find slightly more asymmetry in the flows here than in the synchronous case, asymmetry which we attribute to the time-dependent radiative forcing profile at least slightly disrupting the $m=1$ standing wave structure. Other evidence for the time-dependent radiative forcing can be found in the temperature profile of the non-synchronous mean night-side, which is a hint hotter here than in an equivalent synchronous model, likely as a result of remnant thermal energy (that is still dissipating) from when that physical region of the atmosphere was irradiated.
On the other hand, when the difference between the rotational and orbital periods is large, other than in the fully radiatively controlled regions of the very outer atmosphere, significant changes to the wind and temperature structure become evident, changes which we can be seen in Figure 16.
At low pressure, e.g. $0.0026$\mathrm{bar}$$, we find that the ‘mean’ (where the mean is taken in the frame of reference of the irradiation) hot-spot lies slightly to the west, approximately $15^{\circ}$, of the sub-stellar point. Such a westwards shift, if paired with opacity sources such as clouds which lead to phase curves probing low-pressure regions of the atmospheres, may explain the significant westwards phase curves observed for CoRoT-2b (Dang et al. 2018). Note that a westward shift in the planetary-scale wave pattern at similarly low pressures was proposed by 2022ApJ...941..171L as an explanation for westward hot-spot offsets. Moving deeper, at $0.016$\mathrm{bar}$$, we find strong thermal ‘wings’ that extend from mid-latitudes on the ‘mean’ day-side to near the equator on the ‘mean’ night-side, ‘wings’ which trace the relatively weak off-equator standing wave circulation the develops here. These ‘wings’ only increase in strength as we move deeper into the atmosphere ($0.2$\mathrm{bar}$$ 16(c)), where we find that the non-synchronous mean day-side temperature map has becoming increasingly smeared eastwards of the (moving) sub-stellar point. Whilst this description make the temperature map sound a little like the classical thermal butterfly, differences in advective transport mean that this is not the case.
@@
Note that the dynamics are slightly different when the rotational period is less than the orbital period. This occurs for two reasons, the first is that the direction of the hot-spot shift is reversed westwards rather than eastwards and the second in the decreasing influence of rotation at low $\Omega_{\mathrm{rot}}$ resulting in divergent rather than rotational dynamics. Note that the aforementioned smearing of the temperature map by the shifting sub-stellar point, as well as the above factors, means that we also no longer find any evidence for the formation of the night-side hot-spot that was a defining feature of our slow-rotation-regime models (subsection 3.3).
@@
(a) $R_{0}$ vs. $U\right>_{equator}$
(b) $R_{0}$ vs. $W\right>_{pole}$
Figure 18: Global mean vertical enthalpy flux for 13 non-synchronous HD209458b-like models spanning the rotation range $\Omega_{\mathrm{rot}}=0.5\Omega_{\mathrm{orb}}\rightarrow 4\Omega_{\mathrm{orb}}$, where $\Omega_{\mathrm{orb}}=\Omega_{0}$ is held constant. Note that negative fluxes indicate net downward enthalpy transport. Note that there is a slight scatter in the mean vertical enthalpy flux around $\Omega_{\mathrm{rot}}=\Omega_{\mathrm{orb}}=\Omega_{0}$ can be linked with the averaging window and the current state on long-timescale oscillations in the atmosphere. The main takeaway from this figure should be that the vertical enthalpy flux rapidly drops off when the different between the orbital and rotational periods increases.
@@ 5 Discussion and Conclusion
This work builds upon previous studies of the influence of orbtial radii on atmospheric dynamics, such as Showman et al. (2009, 2015), by isolating the effects of rotation from the change in insolation that occurs when modifying the orbital radii. Our analysis revealed that, for the rotation rate ranges considered, our models fall into one of a number of distinct regimes.
For synchronous rotation we find three distinct dynamical regimes: slow rotation when $\Omega_{\mathrm{rot}}<0.5\Omega_{0}$, the classical hot Jupiter regime when $\sim 0.5\Omega_{0}<\Omega_{\mathrm{rot}}<\sim 5\Omega_{0}$ (an analysis of which can be found in the main text), and the rapidly rotating regime when $\Omega_{\mathrm{rot}}\gg 5\Omega_{0}$. Note that, as discussed in subsection 3.6, these regimes are broadly in line with the rotation regimes for terrestrial exoplanet atmosphere dynamics discussed by Haqq-Misra et al. (2018), albeit with differences in dynamics due to differences between the insolation of hot Jupiters and habitable zonal terrestrial exoplanets. They are also different from the rotation regimes discussed by Showman et al. (2015) in that we do not consider the weakly irradiated regime that they find at larger orbital radii and our rapidly rotating regime maintains a strong zonal jet due to the combination of significant, synchronous, day-side irradiation (i.e. a high day-night temperature contrast) and the suppression of off-equator dynamics by the coriolis effect in our models (we consider much shorter orbital periods than Showman et al. 2015).
@@
We find that the slowly rotating regime is characterised by a slower zonal-mean zonal-jet (1(a)) which does not extend deep into the atmosphere ($P>0.5$\mathrm{bar}$$ is quiescent). The underlying cause of this slow wind can be understood via the Helmholtz wind decomposition, which reveals that the strongest component of the horizontal wind, in the outer atmosphere, is the divergent component associated with a global overturning circulation as heat rises on the day-side, diverges from the sub-stellar point, before converging on the night-side, just to the west of the sub-stellar point. Note that the exact location at which the winds converge is dependent upon the rotation rate, with the convergence point shifting westwards towards the eastern terminator as the influence of rotation grows. This divergent wind, and the associated global overturning circulation acts as the main driver of vertical enthalpy transport and deep adiabatic heating, leading to the formation of a rather unexpected feature: a night-side hot-spot (5(g), 10(a), and 9(b)). This night-side hot-spot appears to be a robust feature of the slowly rotating regime and plays a role in shaping the mid-atmosphere dynamics. For example, around $0.2$\mathrm{bar}$$, we find that the night-side hot spot is hotter than any point on the day-side, an effect which reshapes the winds, driving a strong night-day flow.
@@
As a result we find that the aforementioned global overturning circulation, associated with the divergent day-side wind and a broad sub-stellar point upflow, is closed in the mid-atmosphere, helping to explain the quiescent deep atmosphere. This effect can be seen in the meridional circulation profile, where we find an equatorial downflow that is confined to low pressures, and horizontal mixing at the bottom of the jet, where the night-side hot-spot forms. Finally this lack of a strong equatorial downflow that connects the outer atmosphere with the deep, $P>10\,$\mathrm{bar}$$, atmosphere means that the net deep heating is also weak (11(a) and 11(d)), leading to a deep atmosphere that is not significantly heated (10(a)), thus suggesting that slowly rotating warm/hot Jupiters should exhibit little to no radius inflation.
@@
As for where we might observe such a planet, including a the possible presence of a night-side hot-spot, the most likely scenario is in a longer orbit around a K/M dwarf. Here the synchronisation time-scale is no longer an issue due to the expected lifetimes of cool stars (greater than the current age of the universe), although the reduced insolation may effect the strength of the night-side heating.
@@
On the other hand, we find that the rapidly rotating regime is characterised by a narrow, but deep, equatorial jet braced by easterly counterflows (1(c)). This is a result of the increasing influence that the Coriolis effect plays on the atmospheric dynamics, suppressing high-latitude off-equator flows. Specifically we find that, as the rotation rate rises, the standing $m=1$ Rossby and Kelvin waves which drive the archetypal super-rotating jet become increasingly tilted and confined to low latitudes, leading not only to the narrower jet seen here as the circulation cells become latitudinally compressed, but also a weakening of the overall equatorial winds as the angular momentum pumping (from mid-latitudes to the equator) mechanism becomes less efficient. Note that this compression only grows as we increase the rotation rate, leading to noticeably slow jets at very high $\Omega_{\mathrm{rot}}$ (17(a)). This latitudinal compression of the zonal-winds and off-equator flows also impacts the advection of enthalpy away from the sub-stellar point. For example, we find that the thermal butterfly (5(c)/h/l) has become somewhat confined to equatorial regions, with eastwards advection at the equator, westwards/polewards advection at low latitudes and almost no advection at higher latitudes where the Coriolis forces are strongest (and winds are weakest). Note that this slight westwards advection at low latitudes might explain the weak westwards phase-curve offsets found in observations by, for example, May et al. (2022).
@@
Vertically, the latitudinal compression and resulting banding of the zonal-mean zonal-wind leads to a meridional circulation profile (4(c)) which contains multiple circulation cells in each hemisphere, although these circulations cells do still combine at the equator to drive a weakened downflow at all pressures, connecting the outer and deep atmospheres. Note that, much like the zonal wind, the meridional circulation is significantly weakened at high latitudes thanks to the Coriolis effect, hence most vertical mixing is also confined to low-latitudes (i.e. the equatorial region). This has important implications for the vertical enthalpy transport and hence heating of the deep atmosphere, both of which we find to be globally weak (8(g), 11(f), and 10(c)). Note however that this is just the global value. If we instead consider the equatorial mean (where vertical mixing is strongest), we find a net downward enthalpy flux comparable to that found in the classical hot Jupiter regime, and which leads to slight warming of the equatorial regions deep T-P profile (and only the equatorial regions due to the quiescence of the off-equator deep atmospheric winds). As such, we might expect such objects to be slightly puffed up near the equator, although such an effect is likely to remain minor and may be entangled with other effects that rotation and irradiation has on the atmospheres shape.
@@
This may help to explain the radius inflation, or lack-thereof, observed for ultra-hot Jupiters: planets with an orbital period of less than around 20 hours tend to exhibit little to no radius inflation e.g. Wasp-18b - Arcangeli et al. 2019 - or Wasp-43b - Hellier et al. 2011), whereas planets with orbital periods of $>22$ hours start to show signs of inflation (e.g. Wasp-103b - Gillon et al. 2014 - or KELT-16b - Oberst et al. 2017), similar to the transition seen with rotation rate here. This is also in agreement with the results of Sainsbury-Martinez et al. (2021), who found similar dynamics and a similar lack of radius inflation for SDSS1411b. Note that these examples suggest a rather sharp transition, however when we consider the larger sample of known Jupiter-sized exoplanets, the transition is less distinct. As shown in Figure 19, which plots the relationship between planetary Radius, Effective Temperature, and Orbital Periods, as expected there appears to be a general relationship between increased effective temperature (i.e. insolation) and planetary radius (Laughlin et al. 2011; Demory & Seager 2011). However, there is also a trend of decreasing radius with decreased orbital periods, albeit with some exceptions/outliers, suggesting that rotation is may play a role in modulating radius inflation.
Figure 19: A plot showing the relationship between the radius and effective temperature (i.e. insolation) of Jupiter-mass planets, and how the rotation rate (represented by the color of each point) modulates inflation. This plot was made using data from exoplanet.eu.
The confinement of dynamics to low latitudes also comes with a corresponding decrease in typical dynamical length-scales, such as the Rossby deformation radius or the Rhine length. In our most rapidly rotating models, these length-scales can approach the resolution of our grid, suggesting that our models might be struggling to fully capture some of the dynamics. As such, we suggest that a future study investigate the most rapidly rotating hot Jupiter regime with a significantly increased horizontal resolution in order to fully resolve the dynamics at play.
@@
Thanks to JWST, as well as other next-generation ground and space based telescopes, there is likely to be in influx of long orbital period hot/warm Jupiters in the near future. Not only would these objects likely fall into the slowly rotating regime, it is also increasingly likely that at least some will not be perfectly synchronously rotating. This can occur because the synchronisation timescale increases with orbital radii, approaching the system age for a hot/warm Jupiter orbiting a sun-like star at $0.2$\mathrm{ua}$$. There are also a number of scenarios which might lead to a non-synchronous but highly irradiated hot Jupiter being observed. For example, we know that hot Jupiters cannot have formed so close to their host star and hence must have undergone planetary migration (Lin et al. 1996; Heller 2019). If we where to observe a hot Jupiter that had recently, or is still undergoing, migration, then it might not yet have had time to synchronise. Another possibility is that we may find a hot/warm Jupiter that in stable 3:2 orbital resonance instead of full 1:1 spin synchronisation. This would lead to dynamics that are similar to our slightly non-synchronous models, with the solar insolation ‘slowly’ sweeping across the planetary atmosphere.
@@
As a result, and since many of the dynamics observed in hot Jupiter atmospheres are directly or indirectly linked to the strong day/night thermal forcing (from the day/night divergent flows in the slowly rotating regime to the thermally driven, $m=1$ Kelvin and Rossby standing waves which are responsible for the equatorial jet when rotational effects dominate), we next investigated how varying levels of non-synchronicity, as characterised by a difference between the varying planetary rotation rate and a fixed orbital period affects atmospheric dynamics. Here we found two additional regimes, depending upon the difference between the planetary rotational and orbital periods: when the difference is small, on the order of a few percent ($<15\%$) of the smaller period, we find dynamics that are qualitatively similar to the synchronous regime, just a little weaker. On the other hand when the difference is large, we find significant differences in the dynamics, especially in the driving of the super-rotating jet. This is in agreement with the results of Showman et al. (2009).
@@
Briefly, when the difference between the rotational and orbital periods is small, we find dynamics, including vertical enthalpy transport, that are structurally similar to those found in the synchronous regime, but just a little bit weaker. The biggest difference comes about when we look at the horizontal temperature map in the outer atmosphere, where we find that the ‘sweeping’ of the stellar irradiation essentially acts as a form of day-night transport, with residual heat leading to a warmer night-side and hence a reduced day-night temperature difference.
@@
The above holds true when we also vary $\Omega_{\mathrm{orb}}$. As such, we expect a small difference between the rotational and orbital periods to have little effect on observable features, such as the phase curve (hot-spot location) or radius. Furthermore, since the synchronisation timescale rapidly decreases with decreasing orbital radii it is highly likely that the majority of warm/hot Jupiters detected in the near future will be either synchronously rotating or only weakly non-synchronous. Whilst this is good in one sense, i.e. we can understand the dynamics of such objects quite easily, it also means that, with current tools, there is a distinct possibility that some weakly non-synchronous objects (which are potentially young, depending upon the local synchronisation time-scale), might be erroneously labelled as tidally-locked.
@@
But what about a scenario in which a planet has recently migrated onto a shorter orbit and not yet synchronised its rotation with its host star, leading to a large difference between the rotational and orbital angular rotation rates? Here our models suggest that the dynamics should be distinct from those found in the synchronous regime. For example, we find a zonal-mean zonal-jet (13(b)) which is both slower and shallower than its synchronous counterpart. This has a knock-on effect on the vertical dynamics, weakening both the meridional circulation and vertical enthalpy flux, the latter of which rapidly decreases as the difference between the rotational and orbital periods grows (Figure 18). The underlying cause of these changes can be best understood by exploring the horizontal wind/temperature maps in the outer atmosphere (Figure 16). One immediate observation is that the thermal butterfly that typically forms in the mid-atmosphere of non-slowly-rotating hot Jupiters is not present. Instead, we find that the temperature map has become longitudinally smeared, with the strength of the smearing depending upon $\Delta\Omega$: i.e. as the difference between the rotational and orbital periods increases, the time since the night-side was last irradiated gets smaller, leading to increased residual heating and a hot-spot that appears to trail the irradiation. Not only does this have the effect of reducing the strength of the day/night forcing, it also prevents the formation of the standing-wave structures which drive much of the dynamics of hot Jupiters, including the super-rotating jet and hence meridional circulation and vertical enthalpy flux. This can also shift the transition between the rotational-dominated and divergence-dominated dynamical regimes to faster planetary rotation rates, which will lead to additional changes in the strength of the vertical advection, since divergence-dominated winds are less efficient at vertical transport. Note that the ‘smearing’ of the temperature profile found in the highly non-synchronous regime is similar to non-synchronous models of Showman et al. (2015) - see their Figure 4.
Overall, this suggests that hot/warm Jupiters with large differences between their rotational and orbital periods should be observationally distinct from a synchronously rotating planet at the same orbital radius. This includes both a reduced phase curve amplitude as well as a reduced hot-spot offset, the latter of which could be either to the east or the west of the sub-stellar point depending upon the sign of the difference between the rotational and orbital periods as well as the pressure-level probed. Evidence for a significant ($15^{\circ}$) west-ward hot-spot offset was found at low-pressures (e.g. $P=0.0026$ bar) in our highly non-synchronous model (with $\Omega_{\mathrm{rot}}=4\Omega_{\mathrm{orb}}=4\Omega_{0}$). Such a shift, if paired with atmospheric opacity features, for example clouds, that lead to phase curves observations probing lower pressure regions of the atmosphere, may explain the westwards phase curve offset and broad phase-curve minimum of CoRoT-2b (Dang et al. 2018).
@@
We might also expect to see a reduced level of radius inflation for such objects thanks to the reduced vertical heat flux, although this conclusion is more tenuous since newly formed Jupiter-like objects which have recently undergone planetary migration are likely still dissipating excess energy, associated with for example gravitational collapse, from the formation process, meaning that their envelopes are likely puffed up (Szulágyi et al. 2016). Such a scenario could be confirmed by comparing observed radii with planetary formation models, such as those of Fortney et al. (2007). Note however that migration history can play a role in the observed radius of hot Jupiters when energy is deposited deep in the interior (Mol Lous & Miguel 2020).
@@
In this work we have discussed how rotation alone can significantly impact and alter the dynamics of Hot Jupiters, ranging from the outer atmosphere wind and temperature structure, to the vertical transport of enthalpy. The latter effect may be of particular importance when trying to explain variations in hot Jupiter radii found at fixed irradiation, in particular the lack of inflation observed for some very-short-orbit hot Jupiters and hot Brown Dwarfs.
@@
One interesting discovery which might have significant implications for atmospheric dynamics and chemistry is the formation of a night-side hot-spot (via divergent day-night winds) in the mid-atmosphere of our slowly and synchronously rotating regime models. If such a night-side hot-spot is a robust feature of more slowly rotating hot/warm Jupiters, something that we wish to investigate using a next-generation GCM (such as expeRT/MITgcm) that robustly models radiative heating and cooling at all pressure levels, the associated temperature inversion may significantly alter atmospheric chemistry by acting, for example, as a cold trap and hence reshaping the vertical distribution of key observable models. Understanding such an effect would likely require modelling the atmospheres of slowly rotating hot/warm Jupiters with a GCM that includes both robust radiative transport and dis-equilibrium chemistry, something that is currently far too computationally expensive to robustly investigate. Although work is ongoing to solve this issue; for example, Lee et al. (2023), have coupled a reduced chemical network (‘mini-chem’) to the GCM Exo-FMS, in order to study the effects of non-equilibrium chemistry on the atmospheres of WASP-39b and HD189733b in a computationally efficient manner. As might be expected, this model showed that transport can have a significant affect on the chemistry and local atmospheric composition. However this comes at the cost of calculating a significantly reduced chemical network (12 species without photochemistry) and the need for $\sim 73$ weeks of wall-clock time using 200 processors to model the 200 Earth-years of simulation time considered here.